A farsighted man uses eyeglasses with a refractive power of 1.90 diopters. Wearing the glasses 0.030 m from his eyes, he is able to read books held no closer than 0.275 m from his eyes. He would like a prescription for contact lenses to serve the same purpose. What is the correct contact lens prescription, in diopters?
Answer #1

With Eyeglasses:

Power, P = 1/f = 1.9 D

object distance, u = -(0.275 - 0.03) = -0.245 m

Applying thin lens equation:

1/v - 1/u = 1/f

1/v - 1/(-0.245) = 1.9, solving, we get, v = -0.458 m

therefore, near point of the person is 0.458 + 0.03 = 0.488 m

Now, with contact lens:

object distance, u = - 0.275 m

image distance, v = - 0.488 m

Applying thin lens equation:

1/v - 1/u = 1/f = P

1/(-0.488) - 1/(-0.275) = P

therefore, P = 1.59 Diopters Answer